A police car is moving at 29 m/sand is behind a speeding car moving faster in the same direction as the police car. The radar gun in the police car emits an 8.0x109 Hz electromagnetic wave. The frequency of the returning wave is measured to be 318Hz less than the emitted frequency. Find the speeder's speed.
1
Expert's answer
2012-11-23T03:01:07-0500
z=(318/2) Hz / (8.0*10^9 Hz) = 2*10^(-8)
z=v/c
v=c*z = 3*10^8 m/s * 2 * 10^(-8) = 6 m/s
speedingcar has velocity 6m/s bigger then police car.
Hence, speeding car has velocity 29+6 = 35 m/s